Show that the points (-2,0), (2,3) and (5,-1) are the vertices of a right triangle. And find its area

1
Expert's answer
2022-06-14T11:22:49-0400

"A(-2,0), B(2,3), C(5, -1)"

"\overrightarrow{BA}=\langle-2-2, 0-3\rangle=\langle-4, -3\rangle"

"\overrightarrow{BC}=\langle5-2, -1-3\rangle=\langle3, -4\rangle"

"\overrightarrow{BA}\cdot\overrightarrow{BC}=-4(3)+(-3)(-4)=0"

Then "\overrightarrow{BA}\perp\overrightarrow{BC}," and we see that the triangle "ABC" with vertices "A(-2,0), B(2,3), C(5, -1)" is a right triangle.


"|\overrightarrow{BA}|=\sqrt{(-4)^2+(-3)^2}=5"

"|\overrightarrow{BC}|=\sqrt{(3)^2+(-4)^2}=5"

"Area_{ABC}=\dfrac{1}{2}|\overrightarrow{BA}||\overrightarrow{BA}|"

"=\dfrac{1}{2}(5)(5)=12.5({units}^2)"